This question needs to be solved using any method and rounded to the nearest hundredth if needed. (and it's a quadratic!)

x2 + x − 30 = 0

Answer by Theo(13342)

(x - 5) * (x + 6) = 0

multiply these factors together using the distributive property to get:

x^2 + 6x - 5x - 30 = 0

combine like terms to get:

x^2 + x - 30 = 0

(x - 5) * (x + 6) = 0 is your solution.

that gets you x = 5 or x = -6.

either value of x makes x^2 + x - 30 = 0

when x = 5, x^2 + x - 30 = 0 becomes 25 + 5 - 30 = 0 which becomes 0 = 0 which is true.

when x = -6, x^2 + x - 30 = 0 becomes 36 - 6 + 30 = 0 which becomes 0 = 0 which is true.

your solution is x = 5 or x = -6.